# Reconstruction of Caterpillar Tanglegrams

### v1

@@ Abstract.
A tanglegram consists of two rooted binary trees with the same number of leaves and a perfect matching between the leaves of the trees. Given a size-$n$ tanglegram, i.e., a tanglegram for two trees with $n$ leaves, a multiset of induced size-$(n-1)$ tanglegrams is obtained by deleting a pair of matched leaves in every possible way. Here, we analyze whether a size-$n$ tanglegram is uniquely encoded by this multiset of size-$(n-1)$ tanglegrams. We answer this question affirmatively in the case that at least one of the two trees of the tanglegram is a caterpillar tree.
@@ 1. Introduction
In this paper, we study the tanglegram reconstruction problem originally posed by Stephan Wagner and László Székely in 2017: Given a tanglegram $\mathcal{T}$ of size $n$ (where $n$ is the number of leaves in $L$ or $R$), does the multiset $\mathcal{D}^{m}(\mathcal{T})$ of induced size-$(n-1)$ tanglegrams obtained from $\mathcal{T}$ by deleting a pair of matched leaves from $\mathcal{T}$ in every possible way uniquely determine $\mathcal{T}$ for $n$ sufficiently large? This problem is motivated by the famous graph reconstruction conjecture that states that every graph $G$ on at least three vertices is uniquely determined (up to isomorphism) from the multiset of graphs obtained by deleting one vertex in every possible way from $G$ [8, 16]. This conjecture has been proven for many special classes of graphs but remains open in general (see [2, 7, 14] for an overview). It has also motivated various other reconstruction problems such as for partially ordered sets [10], permutations [15], and groups [12].
@@
The remainder of the paper is organized as follows. In Section 2, we introduce general terminology and notation. We then define different types of trees in Section 3. In Section 4, we consider the tanglegram reconstruction problem involving caterpillar trees and prove Theorem 1.1.
@@ 2. Preliminaries
A rooted tree is a tree $T$ where a special vertex, denoted $r_{T}$, is identified as the root of $T$. The parent of a non-root vertex $v$ is $v$’s neighbor on the $r_{T}-v$ path in $T$ and $y$ is a child of $v$ if $v$ is the parent of $y$. The leaves of $T$ are the vertices with no children. Two leaves with a common parent are called a cherry. We let $\mathbb{L}(T)$ denote the set of leaves and $|T|$ denote the number of leaves. We refer to $|T|$ as the size of $T$.
@@
We remark that whenever we write $T=T^{\prime}$, we mean that $T$ and $T^{\prime}$ are equal up to an isomorphism that maps root to root. In general, a rooted binary tree can be drawn in the plane in multiple ways, as shown in Figure 1
Figure 1. Two drawings for the same rooted binary tree $T$.
@@
Let $T_{1},T_{2}$ be rooted binary trees and let $T_{1}\oplus T_{2}$ be the rooted binary tree $T$ obtained by taking a new vertex $r_{T}$ and joining the roots of $T_{1}$ and $T_{2}$ to $r_{T}$. If $T=T_{1}\oplus T_{2}$, we say $T$ is composed of $T_{1}$ and $T_{2}$. Furthermore, we call $T_{1}$ and $T_{2}$ the maximal pending subtrees of $T$.
Since we consider unlabeled rooted binary trees up to isomorphism, the operation $\oplus$ is commutative. However, if not stated otherwise, for $T=T_{1}\oplus T_{2}$, we assume $|T_{1}|\leq|T_{2}|$.
@@
The caterpillar on $n$ leaves, denoted $C_{n}$, is the rooted binary tree $C_{n}=C_{1}\oplus C_{n-1}$, where $C_{1}$ is the unique rooted binary tree of size 1.
Alternatively, one can view the caterpillar of size $n$ as the unique tree in which the removal of all leaves results in the path graph on $n-1$ vertices, such that the root of the tree is an endvertex of the path. Note that the tree in Figure 1 is $C_{5}$.
We now turn to induced subtrees and decks.
@@
Let $T$ be a rooted binary tree with leaf set $\mathbb{L}(T)$ and let $S\subseteq\mathbb{L}(T)$. The rooted binary tree induced by $S$, denoted $T[S]$, is the tree formed by starting with the minimal subtree of $T$ containing all the vertices in $S$, declaring as root its closest vertex to the root of $T$, and then suppressing all internal vertices with only one child. We also refer to $T[S]$ as an induced subtree of $T$.
@@
The deck $\mathcal{D}(T)$ of $T$ is the set of trees $T^{\prime}$ with $|T^{\prime}|=|T|-1$ such that $T^{\prime}$ is an induced subtree of $T$. The multideck $\mathcal{D}^{m}(T)$ of $T$ is the multiset where the multiplicity of every $T^{\prime}\in\mathcal{D}(T)$ is |$\left|\{S\subseteq\mathbb{L}(T):T^{\prime}=T[S]\}\right|$.
We often refer to the elements of $\mathcal{D}^{m}(T)$ (respectively $\mathcal{D}(T)$) as cards. We say that a size-$n$ tree $T$ is reconstructable from its deck (resp. multideck) if whenever $T_{1}$ is a size-$n$ tree with $T_{1}\neq T$, we have $\mathcal{D}(T)\neq\mathcal{D}(T_{1})$ (resp. $\mathcal{D}^{m}(T)\neq\mathcal{D}^{m}(T_{1})$).
@@
A tanglegram $\mathcal{T}=(L,R,\sigma)$ consists of two rooted binary trees, the left tree $L$ and the right tree $R$, along with a perfect matching $\sigma$ between their leaves. We will denote elements of $\sigma$ using edges $vu$, where $v\in\mathbb{L}(L)$ and $u\in\mathbb{L}(R)$. For an edge $vu\in\sigma$, we also say that $v$ and $u$ are matched. The size of a tanglegram is the common size of its underlying trees.
An isomorphism between two tanglegrams is a graph isomorphism that maps the root of the left (right) tree to the root of the left (right) tree. A tanglegram layout is a straight line drawing in which the two trees are drawn without crossing on the two sides of a strip and the matching edges are drawn inside the strip. Tanglegrams can have several layouts, as shown in Figure 2.
Figure 2. Two drawings of the same tanglegram.
@@
When $E=\sigma-\{uv\}$ for some $vu\in\sigma$, we abuse notation and also denote $\mathcal{T}[\sigma-\{uv\}]$ as $\mathcal{T}-vu$. If the focus is on a particular leaf removed from $L$ (respectively $R$), then we abuse notation and also denote this as as $\mathcal{T}-v$ (respectively $\mathcal{T}-u$).
@@
The deck $\mathcal{D}(\mathcal{T})$ of the tanglegram $\mathcal{T}$ is the set of size $|\mathcal{T}|-1$ tanglegrams that are induced subtanglegrams of $\mathcal{T}$. The multideck $\mathcal{D}^{m}(\mathcal{T})$ of the tanglegram $\mathcal{T}$ contains each element $\mathcal{T}^{\prime}\in\mathcal{D}(\mathcal{T})$ with multiplicity |$|\{E\subseteq\sigma:\mathcal{T}^{\prime}=\mathcal{T}[E]\}|$.
As for trees, we often refer to the elements of $\mathcal{D}^{m}(\mathcal{T})$ (respectively $\mathcal{D}(\mathcal{T})$) as cards. An example is shown in Figure 3.
$\mathcal{T}$
--$\mathcal{T}_{1}$--$\mathcal{T}_{2}$--$\mathcal{T}_{3}$
@@ 3. Types of trees
In this section, we partition the set of rooted binary trees into three types. This partition will be used in the next section to prove our main theorem.
@@
Let $i\in\mathbb{N}$ and $T_{1},T_{2}$ be rooted binary trees. We define the operations $(T_{1}\oplus)^{i}T_{2}$ as follows: $(T_{1}\oplus)^{0}T_{2}=T_{2}$ and for $i>0$, $(T_{1}\oplus)^{i}(T_{2})=T_{1}\oplus((T_{1}\oplus)^{i-1}T_{2})$.
@@
Let $T$ be a rooted binary tree. Then $T$ is strippable if $|T|\geq 2$ and $T=C_{1}\oplus T^{\prime}$ for some tree $T^{\prime}$. Otherwise, $T$ is non-strippable. The stripping of $T$ is $(i,Q)$, where $Q$ is a non-strippable tree, $i$ is a nonnegative integer, and $T=(C_{1}\oplus)^{i}Q$. We call the leaves from the $C_{1}$’s strippable in $T$. Note that $i=0$ indicates that $T=T^{\prime}$, and that $T$ is non-strippable.
Note that $C_{n}$ has stripping $(n-1,C_{1})$. When $T$ is strippable, we will often label the strippable leaves according to their distance from the root (the only time this is ambiguous is the two leaves of $C_{n}$ at distance $n-1$ from the root).
@@
Let $T$ be a strippable tree with stripping $(i,Q)$. A labeling of the strippable leaves with subscripted symbols $\{v_{1},\ldots,v_{i}\}$ is a distance-to-root labeling if $v_{k}$ is distance $k$ from $r_{T}$ for $1\leq k\leq i$. When $T=C_{n}$, the distance-to-root labeling gives the remaining vertex label $v_{n}$.
@@
Let $T$ be a rooted binary tree. We say $T$ is type 00 if it is a caterpillar, type $1$ if it is strippable and not a caterpillar, and type $2$ otherwis
Notice that if $T$ is type $1$ with stripping $(i,Q)$, then $|Q|\geq 4$ and $|T|=|Q|+i\geq 5$. If $T$ is type $2$, then $T=R_{1}\oplus R_{2}$ for some rooted binary trees $R_{1},R_{2}$ with $|R_{1}|,|R_{2}|\geq 2$.
In Figure 4, we depict the three rooted binary trees of size 5. They are of types 0, 1, and 2, respectively.
Figure 4. The three trees of size $5$, respectively of types 0, 1, and 2.
4. Reconstruction of caterpillar tanglegrams
In this section, we prove reconstruction for tanglegrams where at least one tree is a caterpillar.
Throughout, we freely use the fact that if $\mathcal{T}=(L,R,\sigma)$ is a tanglegram of size $n\geq 5$, the multideck $\mathcal{D}^{m}(\mathcal{T})$ uniquely determines the trees $L$ and $R$. This is a direct consequence of Theorem 2.7 and is formalized by the following corollary.
Corollary 4.1.
For any tanglegram $\mathcal{T}=(L,R,\sigma)$ of size $n\geq 5$, the trees $L$ and $R$ are reconstructable from the multideck $\mathcal{D}^{m}(\mathcal{T})$.
We start with the case when both trees are caterpillars and use the distance-to-root labelings $\{v_{1},\ldots,v_{n}\}$ and $\{u_{1},\ldots,u_{n}\}$ of $L$ and $R$, respectively. Furthermore, for a card $(L^{\prime},R^{\prime},\sigma^{\prime})\in\mathcal{D}^{m}(\mathcal{T})$, we use subscripted symbols $\{v_{1}^{\prime},\ldots,v_{n-1}^{\prime}\}$ (resp. $\{u_{1}^{\prime},\ldots,u_{n-1}^{\prime}\}$), for the distance-to-root labeling of the induced subtree $L^{\prime}$ (resp. $R^{\prime}$).
We first handle the case in which $v_{1}$ is matched to the cherry of $R$.

### v2

@@ Abstract.
A tanglegram consists of two rooted binary trees with the same number of leaves and a perfect matching between the leaves of the trees. Given a size-$n$ tanglegram, i.e., a tanglegram for two trees with $n$ leaves, a multiset of induced size-(n1)(n-1) tanglegrams is obtained by deleting a pair of matched leaves in every possible way. Here, we analyze whether a size-$n$ tanglegram is uniquely encoded by this multiset of size-(n1)(n-1) tanglegrams. We answer this question affirmatively in the case that at least one of the two trees of the tanglegram is a caterpillar tree.
@@ 1. Introduction
In this paper, we study the tanglegram reconstruction problem originally posed by Stephan Wagner and László Székely in 2017: Given a tanglegram $\mathcal{T}$ of size $n$ (where $n$ is the number of leaves in $L$ or $R$), does the multiset $\mathcal{D}^{m}(\mathcal{T})$ of induced size-(n1)(n-1) tanglegrams obtained from $\mathcal{T}$ by deleting a pair of matched leaves from $\mathcal{T}$ in every possible way uniquely determine $\mathcal{T}$ for $n$ sufficiently large? This problem is motivated by the famous graph reconstruction conjecture that states that every graph $G$ on at least three vertices is uniquely determined (up to isomorphism) from the multiset of graphs obtained by deleting one vertex in every possible way from $G$ [10, 19]. This conjecture has been proven for many special classes of graphs but remains open in general (see [2, 9, 17] for an overview). It has also motivated various other reconstruction problems such as for partially ordered sets [12], permutations [18], and groups [15].
@@
The remainder of the paper is organized as follows. In Section 2, we introduce general terminology and notation. We then define different types of trees in Section 3. In Section 4 we discuss tanglegrams of size at most $5$ – in particular we explain that Theorem 1.1 is best possible in the sense that the lower bound $n\geq 6$ can not be changed to $n\geq 5$. In Sections 57, we consider the tanglegram reconstruction problem involving caterpillar trees, distinguishing three cases, depending on the type of the tree the caterpillar is matched with. In Section 8 we finally prove Theorem 1.1 and pose a remaining problem.
@@ 2. Preliminaries
A rooted tree is a tree $T$ where a special vertex, denoted $r_{T}$, is identified as the root of $T$. The parent of a non-root vertex $v$ is $v$’s neighbor on the $r_{T}-v$ path in $T$ and $y$ is a child of $v$ if $v$ is the parent of $y$. The leaves of $T$ are the vertices with no children. Two leaves with a common parent are called a cherry. We let $\mathbb{L}(T)$ denote the set of leaves and $\left\langle T\right\rangle$ denote the number of leaves of $T$. We refer to $\left\langle T\right\rangle$ as the size of $T$.
@@
We remark that whenever we write $T=T^{\prime}$, we mean that $T$ and $T^{\prime}$ are equal up to an isomorphism that maps root to root.
@@
If $T$ is a rooted binary tree, a standard drawing of $T$ is a plane drawing of $T$ such that its leaves lie on a straight line (called the leaf-line of $T$), the tree is drawn with straight lines on one of the two half-planes the line defines in the plane, and if $x,y$ are vertices of $T$ such that $x$ is the parent of $y$, then the distance of $x$ from the leaf-line is larger than the distance of $y$ from the leaf-line.
In general, a rooted binary tree has multiple standard drawings, as shown in Figure 1.
Figure 1. Two standard drawings for a rooted binary tree.
@@
Let $T_{1},T_{2}$ be rooted binary trees and let $T_{1}\oplus T_{2}$ be the rooted binary tree $T$ obtained by taking a new vertex $r_{T}$ and joining the roots of $T_{1}$ and $T_{2}$ to $r_{T}$. If $T=T_{1}\oplus T_{2}$, we say $T$ is composed of $T_{1}$ and $T_{2}$. Furthermore, we call $T_{1}$ and $T_{2}$ the pending subtrees of $T$.
The operation $\oplus$ is commutative. However, if not stated otherwise, for $T=T_{1}\oplus T_{2}$, we assume $\left\langle T_{1}\right\rangle\leq\left\langle T_{2}\right\rangle$.
@@
The caterpillar on $n$ leaves, denoted $C_{n}$, is the rooted binary tree defined recursively as follows: $C_{1}$ is the unique rooted binary tree of size 1, and for $n>1$, $C_{n}=C_{1}\oplus C_{n-1}$.
Alternatively, one can view the caterpillar of size $n>1$ as the unique tree in which the removal of all leaves results in the path graph on n1n-1 vertices, such that the root of the tree is an endvertex of the path. Note that the tree in Fig. 1 is $C_{5}$.
We next turn to induced subtrees and decks. We first recall an operation on graphs. Let $G$ be a graph and let $w$ be a degree-2 vertex with neighbors $u$ and $v$. Then, suppressing $w$ means deleting $w$ and its two incident edges, $uw$ and $vw$, and introducing the edge $uv$. Note that this is the same as contracting either of the edges $uw$ or $vw$.
@@
Let $T$ be a rooted binary tree with leaf set $\mathbb{L}(T)$ and let $S\subseteq\mathbb{L}(T)$. The rooted binary tree induced by $S$, denoted $T[S]$, is the tree formed by starting with the minimal subtree of $T$ containing all the vertices in $S$, declaring as root its closest vertex to the root of $T$, and then suppressing all non-root degree-2 vertices. We also refer to $T[S]$ as an induced subtree of $T$.
Definition 2.7.
The deck $\mathcal{D}(T)$ of $T$ is the set of trees $T^{\prime}$ with $\left\langle T^{\prime}\right\rangle=\left\langle T\right\rangle-1$ such that $T^{\prime}$ is an induced subtree of $T$. The multideck $\mathcal{D}^{m}(T)$ is a multiset with the same elements as the deck, but they come with multiplicities. To determine these multiplicities, label the leaves of $T$ arbitrarily, and denote the set of labeled leaves by $\mathbb{L}^{*}(T)$. The multiplicity of an element $T^{\prime}\in\mathcal{D}(T)$ in $\mathcal{D}^{m}(T)$ is $\left|\{S\subseteq\mathbb{L}^{*}(T):T^{\prime}=T[S]\}\right|$.
We often refer to the elements of $\mathcal{D}(T)$ (respectively $\mathcal{D}^{m}(T)$) as cards.
Definition 2.8.
A tree $T$ is reconstructable from its deck (resp. multideck) if for any tree $T^{\prime}$, $\mathcal{D}(T^{\prime})=\mathcal{D}(T)$ (resp. $\mathcal{D}^{m}(T^{\prime})=\mathcal{D}^{m}(T)$) implies $T=T^{\prime}$.
@@
A tanglegram $\mathcal{T}=(L,R,\sigma)$ consists of two rooted binary trees, the left tree $L$ and the right tree $R$, along with a perfect matching $\sigma$ between their leaves. We will denote elements of $\sigma$ using edges $vu$, where $v\in\mathbb{L}(L)$ and $u\in\mathbb{L}(R)$. For an edge $vu\in\sigma$, we also say that $v$ and $u$ are matched. The size $\left\langle\mathcal{T}\right\rangle$ of a tanglegram $\mathcal{T}$ is the common size of its underlying trees, $|\sigma|$.
An isomorphism between two tanglegrams is a graph isomorphism that maps the root of the left (right) tree to the root of the right (left) tree.
Definition 2.11.
A layout of a tanglegram $\mathcal{T}=(L,R,\sigma)$ is a straight line drawing such that for some $a<b$ we have a standard drawing of $L$ with a vertical leaf-line $x=a$ to the left of the leaf-line, a standard drawing of $R$ with a vertical leaf-line $x=b$ to the right of the leaf-line, and the edges in $\sigma$ are drawn as straight lines.
Tanglegrams can have several layouts, as shown in Figure 2.
Figure 2. Two layouts of the same tanglegram.
@@
When $E=\sigma-\{vu\}$ for some $vu\in\sigma$, we abuse notation and also denote $\mathcal{T}[\sigma-\{vu\}]$ as $\mathcal{T}-vu$. If the focus is on a particular leaf removed from $L$ (respectively $R$), then we abuse notation and also denote this as as $\mathcal{T}-v$ (respectively $\mathcal{T}-u$).
@@
The deck $\mathcal{D}(\mathcal{T})$ of the tanglegram $\mathcal{T}$ is the set of size $\left\langle\mathcal{T}\right\rangle-1$ tanglegrams that are induced subtanglegrams of $\mathcal{T}$. The multideck $\mathcal{D}^{m}(\mathcal{T})$ of the tanglegram $\mathcal{T}$ is a multiset whose elements are the elements of $\mathcal{D}(\mathcal{T})$. To determine these multiplicities, label the edges of $\sigma$ arbitrarily, and denote the labeled set by $\sigma^{*}$. The multiplicity of an element $\mathcal{T}^{\prime}\in\mathcal{D}(\mathcal{T})$ in $\mathcal{D}^{m}(\mathcal{T})$ is $|\{E\subseteq\sigma^{*}:\mathcal{T}^{\prime}=\mathcal{T}[E]\}|$.
As for trees, we often refer to the elements of $\mathcal{D}^{m}(\mathcal{T})$ (respectively $\mathcal{D}(\mathcal{T})$) as cards. An example is shown in Fig. 3.
$e_{1}$$e_{2}$$e_{3}$$e_{4}$$e_{5}$$\mathcal{T}$
--$\mathcal{T}_{1}=\mathcal{T}-e_{1}=\mathcal{T}-e_{2}$--$\mathcal{T}_{2}=\mathcal{T}-e_{3}$--$\mathcal{T}_{3}=\mathcal{T}-e_{4}=\mathcal{T}-e_{5}$
@@
Definition 2.14.
A tanglegran $\mathcal{T}$ is reconstructable from its deck (resp. multideck) if for any tanglegram $\mathcal{T}^{\prime}$, $\mathcal{D}(\mathcal{T}^{\prime})=\mathcal{D}(\mathcal{T})$ (resp. $\mathcal{D}^{m}(\mathcal{T}^{\prime})=\mathcal{D}^{m}(\mathcal{T})$) implies $\mathcal{T}=\mathcal{T}^{\prime}$.
We are interested in determining whether all tanglegrams or all tanglegrams of a certain class can be reconstructed from their deck. This brings up the following problem: Say, we have a tanglegram class $\mathcal{H}$ (e.g. $\mathcal{H}$ is the class of tanglegrams with at least one of the left or right tree being a caterpillar) and we know that all tanglegrams in $\mathcal{H}$ have unique decks, but there are some decks that also correspond to tanglegrams not in $\mathcal{H}$. In this case, given a deck of a tanglegram, we still have no information on whether we can reconstruct the tanglegram. Because of this, we require in our definition that we can determine from the deck (or multideck) whether the tanglegram it belongs to must be in $\mathcal{H}$ or not.
Definition 2.15.
Let $n$ be a positive integer, and $\mathcal{D}$ be a set (resp. multiset) of size n1n-1 tanglegrams. We denote by $\mathbb{T}[\mathcal{D}]$ the set of size $n$ tanglegrams with deck (resp. multideck) $\mathcal{D}$. For tanglegram class $\mathcal{H}$ and positive integer $n$, we say that $\mathcal{H}$ is deck-decidable (resp. multideck-decidable) for size $n$ if for every set $\mathcal{D}$ (resp. multiset) of size n1n-1 tanglegrams with $\mathbb{T}(\mathcal{D})\neq\emptyset$, either $\mathbb{T}[\mathcal{D})]\subseteq\mathcal{H}$ or $\mathbb{T}[\mathcal{D}]\cap\mathcal{H}=\emptyset$, and we can decide from $\mathcal{D}$ which is the case.
Definition 2.16.
Let $\mathcal{H}$ be a tanglegram class and $n$ be a positive integer. The tanglegrams of $\mathcal{H}$ of size $n$ are reconstructable from their deck (resp. multideck) if $\mathcal{H}$ is deck-decidable (resp. multideck-decidable) for size $n$ and for every $\mathcal{T}\in\mathcal{H}$ of size $n$, $\mathcal{T}$ is reconstructable from its deck (resp. multideck).
Note that $\mathcal{D}(\mathcal{T})$ (respectively $\mathcal{D}^{m}(\mathcal{T})$) includes all elements of $\mathcal{D}(L)$ and $\mathcal{D}(R)$ (respectively $\mathcal{D}^{m}(L)$ and $\mathcal{D}^{m}(R)$). The following lemma is therefore immediate from Theorem 2.9.
Lemma 2.17.
Let $\mathcal{T}=(L,R,\sigma)$ be a tanglegram. If $\left\langle\mathcal{T}\right\rangle\geq 6$, then $L$ and $R$ can be reconstructed from $\mathcal{D}(\mathcal{T})$, and if $\left\langle\mathcal{T}\right\rangle=5$, then $L$ and $R$ can be reconstructed from $\mathcal{D}^{m}(\mathcal{T})$. Consequently, the tanglegram class containing all tanglegrams where at least one of the left or right tanglegram is a caterpillar is deck-decidable for size at least $6$.
Thus, our algorithmic approach reduces the tanglegram reconstruction problem to reconstructing the set $\sigma$ of matching edges from $\mathcal{D}^{m}(\mathcal{T})$.
@@ 3. Types of trees
In this section, we partition the set of rooted binary trees into three types. This partition will be used in later sections to prove our main theorem.
@@
Let $i\in\mathbb{N}$ and $T_{1},T_{2}$ be rooted binary trees. We define the operation $(T_{1}\oplus)^{i}T_{2}$ as follows: $(T_{1}\oplus)^{0}T_{2}=T_{2}$ and for $i>0$, $(T_{1}\oplus)^{i}(T_{2})=T_{1}\oplus((T_{1}\oplus)^{i-1}T_{2})$.
@@
Let $T$ be a rooted binary tree. Then $T$ is strippable if $T=C_{1}$ or $T=C_{1}\oplus T^{\prime}$ for some tree $T^{\prime}$. Otherwise, $T$ is non-strippable. The stripping of $T$ is defined as follows: If $T=C_{n}$, its stripping is $(n,\emptyset)$. Otherwise, the stripping of $T$ is $(i,Q)$, where $Q$ is a non-strippable tree, $i$ is a nonnegative integer, and $T=(C_{1}\oplus)^{i}Q$. If $T=C_{n}$, all of its leaves are strippable, otherwise, the leaves not in $Q$ are strippable in $T$. Note that $i=0$ indicates that $T=T^{\prime}$, and that $T$ is non-strippable (see, e.g., Fig. 4).
Note that if $(i,Q)$ is a stripping of a tree $T$, then $\left\langle T\right\rangle=i+\left\langle Q\right\rangle$ (where we use $\left\langle\emptyset\right\rangle=0$). Moreover, if $i\geq 1$, then one of the following holds: $i=n$ and $T=C_{n}$; or $\left\langle Q\right\rangle\geq 4$. When $T$ is strippable, we will often label the strippable leaves according to their distance from the root (the only time this is ambiguous is when $T=C_{n}$).
@@
Let $T$ be a strippable tree with stripping $(i,Q)$. A labeling of the strippable leaves with subscripted symbols $\{v_{1},\ldots,v_{i}\}$ is a distance-to-root labeling if for all $k:1\leq k\leq\min(i,n-2)$, $v_{k}$ is distance $k$ from $r_{T}$ for $1\leq k\leq i$. When i>n2i>n-2, then $T=C_{n}$, and the distance-to-root labeling gives the remaining two vertices the labels $v_{n}$ and $v_{n-1}$ (note that these two vertices form a cherry).
@@
Let $T$ be a rooted binary tree. We say $T$ is type 0 if it is a caterpillar, type 1 if it is strippable and not a caterpillar, and type 2 otherwise. Moreover, a type 2 tree is type 2a if it is of the form $Q\oplus Q$ for some rooted binary tree $Q$, type $2b$ if it is of the form $Q\oplus Q^{\prime}$ for some rooted binary tree $Q$ and $Q^{\prime}\in\mathcal{D}(Q)$, and type 2c otherwise.
Notice that if $T$ is type 1 with stripping $(i,Q)$, then $\left\langle Q\right\rangle\geq 4$ and $\left\langle T\right\rangle=\left\langle Q\right\rangle+i\geq 5$. If $T$ is type 2, then $T=Q_{1}\oplus Q_{2}$ for some rooted binary trees $Q_{1},Q_{2}$ with $\left\langle Q_{1}\right\rangle,\left\langle Q_{2}\right\rangle\geq 2$.
In Fig. 4, we depict the three rooted binary trees of size 5. They are of types 0, 1, and 2b, respectively.
$C_{5}$, stripping $(5,\emptyset)$.Tree with stripping $(1,C_{2}\oplus C_{2})$.TT, stripping $(0,T)$.
Figure 4. The three trees of size $5$, respectively of types 0, 1, and 2b.
The following statements are obvious and are implicit in [4]: